# Superconducting Diode Effect in Weak Localization Regime

### v1

@@ Abstract
$\alpha_{4}$ $=-4\pi T\!\sum_{n>0}\sum_{\lambda=\pm}\Big{(}f_{3n\lambda}-\!\frac{1}{6}[f_{1n\lambda}]^{3}\Big{)}.$ (17)
Note that $f_{jn\lambda}$ is derived from Eq. (15), and $\alpha_{2}$ is regularized by introducing the ultraviolet cutoff, which is the Debye energy $\omega_{\rm D}$, to the energy integration. The form of the gap equation in the GL level suggests that the free energy and the supercurrent are given as $f_{\mathrm{GL}}=\frac{1}{2}\alpha_{2}\Delta^{2}_{\rm c}+\frac{1}{4}\alpha_{4}\Delta^{4}_{\rm c}$ and $I_{\rm s}=\partial f_{\mathrm{GL}}/\partial A_{y}$, respectively. The helical SC state for any supercurrent $I_{s}$ is then determined by the latter equation. In particular, the second-order SC phase transition line in the absence of the supercurrent is calculated based on the conditions $\alpha_{2}=0$ and $\partial\alpha_{2}/\partial A_{y}=0$. From now on, let $A_{0}$ be the $A_{y}$ field in the zero-current SC states. We denote the transition temperature at a magnetic field $h$ as $T_{\rm c}(h)$, which is denoted as $T_{\rm c0}(h)$ when we neglect the e-e interaction. At $h=0$, we describe $T_{\rm c0}(0)=T_{\rm c0}$ and $T_{\rm c}(0)=T_{\rm c}$, for simplicity. For later use, we define the reduced temperature $\tau_{\rm r}$ as $\tau_{\rm r}=T/T_{\rm c0}$ or $\tau_{\rm r}=T/T_{\rm c}$.
WL conductivity: When the current reaches its critical value, the SC state changes to a resistive state. Here, we study the WL correction to the conductivity, $\sigma_{\rm WL}$ [37, 38, 39]. For simplicity, we neglect the e-e interactions. We find that it is convenient to work with the alternative normal saddle point $\bar{Q}^{\prime}_{\rm N}=U^{\prime}\Lambda^{\prime}_{\rm N}U^{\prime-1}$, where $\Lambda^{\prime}_{\rm N}=\mathrm{diag}\,\{\gamma_{3},\gamma_{3}\}$ and $U^{\prime}=\mathrm{diag}\,\{\check{U}_{\rm K},\check{U}^{\mathbb{T}}_{\rm K}\}$. Here, $U^{\mathbb{T}}$ denotes the transpose of a matrix and the interchange of the time of $U$, which means that $U^{\mathbb{T}}_{t_{1}t_{2}}=U^{\rm t}_{t_{2}t_{1}}$ with $U^{\rm t}_{t_{2}t_{1}}$ being the ordinary transpose in matrix elements of $U_{t_{2}t_{1}}$. The matrix $\check{U}_{\rm K}$ has the same form as Eq. (4), but $\tau_{0}$ is replaced by $\sigma_{0}$.
This saddle point comes from the time-reversal symmetry condition of the matrix field $Q$. To the best of our knowledge, this saddle point was first introduced in Ref. 40 to calculate the WL conductivity $\sigma_{\rm WL}$ in the spinless case within the Keldysh functional formalism. Here, we generalize their formalism to include spin degrees of freedom. We explain the procedure in [35], and the equation for $\sigma_{\rm WL}$ is consistent with existing results [41, 42, 44, 43].
@@
Numerical results: In the numerical calculations, we fix the parameters, $\Gamma_{\rm c}=0.2$, $\omega_{\rm D}/T_{\rm c0}=0.1309$, and Fermi energy $E_{\rm F}/T_{\rm c0}=p_{\rm F}^{2}/(2mT_{\rm c0})=10^{3}$, and consider the following two sets of parameters: (I) $\alpha/p_{\rm F}=0.01,~\tau T_{\rm c0}=1.5\times 10^{-2}$ and (II) $\alpha/p_{\rm F}=0.002,~\tau T_{\rm c0}=7.5\times 10^{-3}$, where $\alpha$ is the Rashba SOC constant and $\tau$ is the impurity scattering time. Note that the values of $\tau T_{\rm c0}$ in cases (I) and (II) correspond to the characteristic lengths of $l_{\rm F}p_{\rm F}=30$ and 151515, respectively, which are consistent with the WL regime. The characteristic energies of SOC and disorder are represented by $\epsilon_{\rm SOC}=\alpha v_{\rm F}$ and $\epsilon_{\rm D}=1/\tau$, respectively. In case (I), the SOC is comparable with disorder as $\epsilon_{\rm SOC}/\epsilon_{\rm D}=0.6$, while case (II) is in a weak SOC regime as $\epsilon_{\rm SOC}/\epsilon_{\rm D}=0.06$. We show the numerical results for the comparable and weak SOC in Figs. 1 and 2, respectively.
@@
Figure 1: The diode quality factor $\eta$ and the WL correction to conductivity $\sigma_{\rm WL}$ in the intermediate SOC regime ($\alpha/p_{\rm F}=0.01,~\tau T_{\rm c0}=1.5\times 10^{-2}$). The solid (dashed) lines show the results in the interacting (free) case with (without) the e-e interaction. Panels (a) and (c) show the color maps of $\sigma_{\rm WL}$ and $\eta$, respectively. The solid and dashed lines show the second-order superconducting transition line. The bold circles in (a) represent the tricritical points. Panels (b) and (d) show $\sigma_{\rm WL}$ and $\eta$ near the SC transition line, respectively. In these figures, we vary the reduced temperature $\tau_{\rm r}$ and magnetic field $h$ so that the temperature is $T=0.99T_{\rm c}(h)$.
@@
We first show the properties of zero-current SC states. As shown in Figs. 1a and 2a, the e-e interactions suppress the transition temperatures and critical magnetic fields. The gauge field $A_{0}$ corresponding to the momentum of Cooper pairs is also suppressed when the interactions are incorporated [35]. Since there is no contribution from the Cooper channel in the equation determining the second-order SC transition line, the suppression of the transition temperature is attributed to the Coulomb channel. The zero-field transition temperatures are $T_{\rm c}=0.8793T_{\rm c0}$ and $T_{\rm c}=0.6457T_{\rm c0}$ in Figs. 1 and 2, respectively. We have confirmed that $T_{\rm c}$ decreases as the impurity scattering time $\tau$ decreases, while it increases toward $T_{\rm c0}$ as the SOC $\alpha$ increases. The transition lines depicted in Figs. 1a and 2a show the second-order phase transition, and they terminate at the tricritical points at high magnetic fields, indicating the first-order transition, at which $\alpha_{4}$ becomes negative. It is well known that the first-order SC transition occurs in the high field region due to the Pauli paramagnetic effects when we adopt the mean-field theory for spatially uniform $s$-wave states [45]. At magnetic fields higher than the tricritical point, the second-order transition from the normal state to the non-uniform stripe SC state is expected [46]. Since the Cooper channel makes a non-zero contribution to $\alpha_{4}$, both Coulomb and Cooper channels affect the tricritical points.
Figure 2: The WL correction to conductivity $\sigma_{\rm WL}$ in the weak SOC regime ($\alpha/p_{\rm F}=0.002,~\tau T_{\rm c0}=7.5\times 10^{-3}$). Panel (a) shows the color map of $\sigma_{\rm WL}$, which is plotted in the same manner as Fig. 1a. Panel (b) shows $\sigma_{\rm WL}$ near the superconducting transition line in the same manner as Fig. 1b.
Next, let us focus on the quality factor $\eta$ of the SD effect. The diode quality factor $\eta$ is defined as $\eta=(1-\xi)/(1+\xi)$, where $\xi=|I_{\rm sc-}/I_{\rm sc+}|$ and $I_{\rm sc\pm}$ is the critical current in the ±$\pm y$-direction. Because $\eta$ is tiny in the weak SOC regime, we calculate the SDE in the intermediate SOC regime and show the results of $\eta$ in Figs. 1c and 1d. As shown in Fig. 1c, the diode quality factor increases as the system approaches the tricritical point. To show the behaviors of the SDE near the transition line, we plot in Fig. 1d the diode quality factor near the transition line as a function of the reduced temperature. The absolute values of the diode quality factor increase divergently as the tricritical point is approached. We see qualitatively the same behaviors in the absence and presence of the e-e interaction, while the tricritical point is shifted. Although the Coulomb interaction significantly reduces the zero-field transition temperature, the diode quality factor almost coincides in both cases at high reduced temperatures. This behavior is deviated as the tricritical points are approached. Because the tricritical point is shifted by interactions, the absolute values of the diode quality factor in the presence of interactions become smaller than those in the case without interactions.
Finally, we discuss the WL correction to the conductivity $\sigma_{\rm WL}$. The correction is normalized with the Drude conductivity $\sigma_{\rm D}=2\nu D$. As shown in Fig. 1a, interestingly, $\sigma_{\rm WL}$ in the intermediate SOC regime decreases as the temperature and magnetic field increase, indicating stronger localization effects. By examining the plot near the transition lines [Fig. 1b], we conclude that the e-e interaction can suppress the WL correction. However, as shown in Fig. 2a, the qualitatively opposite behaviors manifest themselves in almost the entire region of temperatures and magnetic fields. Looking at the results near the transition lines [Fig. 2b], we find that the e-e interaction can enhance the WL correction when the SOC is weak.
@@
Summary: We have formulated the renormalized GL theory including the Cooper and Coulomb interaction terms from the NLSM based on the Keldysh functional formalism. Our formalism extends the previous one and enables us to study dirty 2D Rashba superconductors under in-plane Zeeman fields, which are typical platforms of the SD effect. We find that the interplay between localization and interactions leads to the corrections that suppress the SC states, the first-order transition, and the SD effect. Furthermore, we have calculated the WL correction to conductivity that characterizes transport when the SC state is destroyed by an electric current exceeding the critical current. The WL correction can be suppressed or enhanced by interactions depending on the strength of SOC. We expect our results to pave the way for switching between superconducting, metallic, and insulating transport via an electric current.

### v2

@@ Abstract
$\alpha_{4}$ $=-4\pi T\!\sum_{n>0}\sum_{\lambda=\pm}\Big(f_{3n\lambda}-\!\frac{1}{6}[f_{1n\lambda}]^{3}\Big),$ (16)
where $\Gamma_{\rm c}$ is the SC coupling constant. Note that $f_{jn\lambda}$ is derived from Eq. (14), and $\alpha_{2}$ is regularized by introducing the ultraviolet cutoff, which is the Debye energy $\omega_{\rm D}$, to the energy integration. Since we are considering helical SC states, the supercurrent $I_{s}$ in the $y$-direction is given as follows [4]:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}I_{\rm s}=\partial f_{\mathrm{GL}}/\partial A_{y}=-\frac{1}{4}\mathcal{C}\frac{\partial}{\partial A_{y}}\,(\alpha^{2}_{2}/\alpha_{4}).$} (17)
The second-order SC phase transition line in the absence of the supercurrent is calculated based on the conditions $\alpha_{2}=0$ and $\partial\alpha_{2}/\partial A_{y}=0$. We denote the transition temperature at a magnetic field $h$ as $T_{\rm c}(h)$, which is denoted as $T_{\rm c0}(h)$ when we neglect the e-e interaction. For simplicity we write $T_{\rm c0}(0)=T_{\rm c0}$ and $T_{\rm c}(0)=T_{\rm c}$. Furthermore, we calculate the critical currents $I_{\rm sc\pm}$ in the $\pm y$-direction, which is obtained by optimizing $I_{\rm s}$, with focus on the diode quality factor defined as $\eta=(1-\xi)/(1+\xi)$ with $\xi=|I_{\rm sc-}/I_{\rm sc+}|$.
@@
In the numerical calculations, we fix the parameters, $\Gamma_{\rm c}=0.2$, $\omega_{\rm D}/E_{\rm F}=0.1309$, and Fermi energy $E_{\rm F}/T_{\rm c0}=p_{\rm F}^{2}/(2mT_{\rm c0})=10^{3}$, where $p_{\rm F}=mv_{\rm F}$ and $m$ is the electron mass. For the Rashba SOC constant $\alpha$ and impurity scattering time $\tau$, we consider the following two sets of parameters: (I) $\alpha/p_{\rm F}=0.01,~\tau T_{\rm c0}=1.5\times 10^{-2}$ and (II) $\alpha/p_{\rm F}=0.002,~\tau T_{\rm c0}=7.5\times 10^{-3}$. Note that the characteristic energies of SOC and disorder are represented by $\epsilon_{\rm SOC}=\alpha v_{\rm F}$ and $\epsilon_{\rm D}=1/\tau$, respectively. Therefore, case (I) is considered a comparable SOC case with $\epsilon_{\rm SOC}/\epsilon_{\rm D}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.3}$, while case (II) is considered a weak SOC case with $\epsilon_{\rm SOC}/\epsilon_{\rm D}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.03}$. Note also that our perturbative approach is valid for large dimensionless conductance $g\gg 1$ [38, 30]. Here, $g$ is given as $g=2\pi\sigma_{\rm D}$ [40], where $\sigma_{\rm D}=2\nu D$ is the Drude conductivity and $\nu=m/(2\pi)$ is the 2D density of states per spin. We have confirmed that the impurity scattering time $\tau$ in both parameter sets (I) and (II) satisfies this condition.
@@
Figure 1: The phase diagram and diode quality factor $\eta$ in the comparable SOC case ($\alpha/p_{\rm F}=0.01,~\tau T_{\rm c0}=1.5\times 10^{-2}$). The solid (dashed) lines show the results in the case with (without) e-e interactions, which is referred to as the interaction (free) case in the main text. (a) The color maps of $\eta$ in the vicinity of the second-order SC transition lines, which change to first-order transitions at the tricritical points represented by the black circles. (c) The transition lines from panel (a) on the normalized temperature-normalized magnetic field scale, namely the $\tau_{\rm r}$-$h_{\rm r}$ scale. (b) and (d) the diode quality factor $\eta$ as a function of $\tau_{\rm r}$ and $h_{\rm r}$, respectively. In these figures, we vary $\tau_{\rm r}$ and $h_{\rm r}$ so that the temperatures satisfy $T=0.99T_{\rm c0}(h)$ in the free case and $T=0.99T_{\rm c}(h)$ in the interaction case. Note that the same $\tau_{\rm r}$ does not imply the same $h$ or $h_{\rm r}$ between the interaction and free cases.
@@
Phase diagram: We first show the properties of zero-current SC states. As shown in Figs. 1a and 3a, the transition temperatures and magnetic fields are lower in the interaction case than in the free case. The gauge field of the helical SC state at zero current is also suppressed when the interactions are incorporated [42]. The zero-field transition temperatures are $T_{\rm c}=0.8793T_{\rm c0}$ and $T_{\rm c}=0.6457T_{\rm c0}$ in Figs. 1a and 3a, respectively. We have confirmed that $T_{\rm c}$ decreases as the impurity scattering time $\tau$ decreases, while it increases toward $T_{\rm c0}$ as the SOC $\alpha$ increases. The second-order transition lines become first-order transitions at the tricritical points, where $\alpha_{4}$ becomes negative. The tricritical temperature and magnetic field in the interaction case, $T^{*}_{\rm c}$ and $h^{*}_{\rm c}$, are lower than those in the free case, $T^{*}_{\rm c0}$ and $h^{*}_{\rm c0}$, respectively. This first-order transition is well known to occur in the high field region due to the Pauli paramagnetic effects [52], and it also occurs when the SOC is sufficiently weak, namely $\Gamma_{\rm r}=2D\alpha^{2}\lesssim 10\Delta_{0}$ where $\Delta_{0}\sim 1.8T_{\rm c0}$ is the SC gap at zero temperature [22, 53]. The parameters used here satisfy this condition. At magnetic fields higher than the tricritical points, although a second-order transition to a stripe SC state is expected in the clean limit [54], it turns out to be a first-order transition to a uniform SC state in the dirty limit [55]. Unlike the previous study [22] that solves the MF gap equation rigorously, our approach cannot describe the first-order transition because the $\Delta_{\rm cl}$-term in the free energy is up to the 4th order.
Figure 2: The diode quality $\eta$ is plotted as a function of the normalized magnetic field $h_{\rm r}$ at normalized temperatures: (a) $\tau_{\rm r}=0.9$ and (b) $\tau_{\rm r}=0.25$. In panel (a), the red and blue squares denote $\eta$ in the interaction and free cases, respectively. The green line shows the fitting by Eq. (18). In panel (b), the red solid and blue dashed lines show $\eta$ in the interaction and free cases, respectively.
SD effect: Let us discuss the SD effect with focus on the diode quality factor $\eta$. Since $\eta$ is tiny in the weak SOC case, we calculate the SD effect in the comparable SOC case and show the color map of $\eta$ in Fig. 1a. It appears that $\eta$ increases as the system approaches the tricritical point. To show the behavior of $\eta$ near the transition line, we plot in Fig. 1b the diode quality factor near the transition line as a function of normalized temperature $\tau_{\rm r}$, defined as $\tau_{\rm r}=T/T_{\rm c0}$ or $\tau_{\rm r}=T/T_{\rm c}$. We see qualitatively the same behavior of $\eta$ in the absence and presence of the e-e interaction. Especially, $\eta$ in the two cases almost coincide at high $\tau_{\rm r}$ despite the significant decrease in the zero-field transition temperature due to the e-e interactions. This means that the SD effect with the same order of $\eta$ as in the free case can be achieved in the interaction case. This finding also suggests a universal law of the diode quality factor, which will be derived below.
In the GL theory [3, 4], the diode quality factor near the second-order transition line follows the scaling law $\eta\sim\mathcal{F}\left[{1-\tilde{\tau}}\right]^{1/2}$. In the free case, the reduced temperature is $\tilde{\tau}=T/T_{\rm c0}(h)$, and the coefficient can be expanded by odd powers of the magnetic field, $\mathcal{F}=\sum_{n=0}A_{2n+1}h^{2n+1}$. In the interaction case, we replace $T_{\rm c0}(h)$ and $A_{2n+1}$ with $T_{\rm c}(h)$ and $A^{\prime}_{2n+1}$, respectively. The relation between $A_{1}$ and $A^{\prime}_{1}$ is derived from the following analysis. We can fit transition lines in Fig. 1a by $T_{\rm c0}(h)/T_{\rm c0}=[1-\chi_{\rm c0}(h/h_{\rm c0})^{2}]^{1/\chi_{\rm c0}}$ and $T_{\rm c}(h)/T_{\rm c}=[1-\chi_{\rm c}(h/h_{\rm c})^{2}]^{1/\chi_{\rm c}}$. These reduce to the known behavior of the transition line $\tau_{\rm r}\simeq 1-h^{2}_{\rm r}$ at low fields, where $h_{\rm r}\equiv h/h_{\rm c0}$ $(h/h_{\rm c})$ and $\tau_{\rm r}=T/T_{\rm c0}$ $(T/T_{\rm c})$. Thus, the transition lines almost coincide on the $\tau_{\rm r}$-$h_{\rm r}$ scale at low fields [Fig. 1c]. We find $h_{\rm c0}=3.61T_{\rm c0}$, $h_{c}=3.28T_{\rm c0}$, $\chi_{\rm c0}=1.89$, and $\chi_{c}=1.84$. Note that $h_{\rm c0}$ and $h_{\rm c}$ are not equivalent to the tricritcal fields ($h^{*}_{\rm c0}=2.6T_{\rm c0}$ and $h^{*}_{\rm c}=2.4T_{\rm c0}$). When the horizontal axis in Fig. 1b is changed to $h_{\rm r}$, the diode quality factors nearly overlap at low fields, as shown in Fig. 1d. This means that $A_{1}/h_{\rm c0}=A^{\prime}_{1}/h_{\rm c}\equiv\mathcal{A}_{1}$, and $\eta$ at low fields follows the universal behavior
$\eta=\mathcal{F}\big[1-\tilde{\tau}\big]^{1/2},\quad\mathcal{F}\simeq\mathcal{A}_{1}h_{\mathrm{r}}.$ (18)
We further find that this law is valid in a wide range of the phase diagram at high temperatures $\tau_{\rm r}>0.9$, as demonstrated in Fig. 2a by setting $\mathcal{A}=2.8\times 10^{-3}\tau^{-1}_{\rm r}$. Based on these analyses, we conclude that the diode quality factor is robust against the corrections due to disorder and e-e interactions in the low-magnetic-field regime.
In previous work [12], a system essentially similar to ours was studied without incorporating e-e interactions. The linear field dependence of $\eta$ is found to be due to the singlet-triplet mixing mechanism. However, unlike our Eq. (18), the analytical expression derived in Ref. [12] cannot describe the whole behavior of $\eta$ in Fig. 2a. Nevertheless, combining their results and Eq. (18) suggests that the interactions may lower both Rashba SOC and transition temperature while preserving the SD effect.
In the low-temperature regime, deviation from the universal law manifests, as shown in Fig. 2b. The roles of the e-e interactions appear non-monotonically. In the high-magnetic-field regime, the diode quality factor $\eta$ is higher in the interaction case than in the free case. This is a result of an increase in the tricritical field in the $h_{\rm r}$ scale due to e-e interactions, and can be qualitatively explained by Eq. (18). However, at low magnetic fields, the diode quality factor is decreased by the interaction.
WL conductivity: When the supercurrent reaches its critical value, the SC state changes to a resistive state. Here, we study the quantum corrections to conductivity in the resistive states. For simplicity, we neglect the corrections to conductivity due to e-e interactions and consider the conventional WL conductivity $\sigma_{\rm WL}$ [44, 45, 46]. The calculation of $\sigma_{\rm WL}$ can be performed in the Keldysh functional formalism by extending the previous work [47] that considers only the spinless case to include the spin degrees of freedom [42]. It appears that our equation for $\sigma_{\rm WL}$ is consistent with existing results [48, 49, 51, 50].
The numerical results of the WL conductivity $\sigma_{\rm WL}$ are shown in Fig. 3. The WL correction normalized by the Drude conductivity is much larger in the weak SOC case [Fig. 3a] than in the comparable SOC case [Fig. 3b]. As the temperature and magnetic field increase, the magnitude of $\sigma_{\rm WL}$ decreases in the weak SOC case but increases in the comparable SOC case. These two results suggest that the localization behaviors are suppressed by the SOC irrespective of Coulomb interactions. In contrast, the magnitude of the diode quality factor $\eta$ with a large SOC is usually larger than with a small SOC for a given temperature and magnetic field. Therefore, our results reveal a trade-off relationship between the SD effect and the WL behavior in the resistive states, at least in the parameter range considered here.
Figure 3: The WL conductivity $\sigma_{\rm WL}$ in (a) the weak SOC regime ($\alpha/p_{\rm F}=0.002,~\tau T_{\rm c0}=7.5\times 10^{-3}$) and (b) the comparable SOC case ($\alpha/p_{\rm F}=0.01,~\tau T_{\rm c0}=1.5\times 10^{-2}$). The solid (dashed) lines show the results in the interaction (free) case. The white circles represent the tricritical points.
@@
Summary: We have formulated the renormalized GL theory including the Cooper and long-range Coulomb interaction terms from the NLSM, based on the Keldysh functional formalism. Our formalism extends the previous one and enables us to study dirty 2D Rashba superconductors under in-plane Zeeman fields, which are typical platforms of the SD effect. We found the robustness of the SD effect with respect to the Cooper and Coulomb interactions near the low-transition magnetic fields, even though the SC states and first-order transition are suppressed by the long-range Coulomb interaction in particular. This means that the magnetic field required to produce a diode quality factor can be reduced by the Coulomb interaction. We also demonstrated a trade-off relationship between the localization behavior in the resistive state, which emerges above the critical current, and the efficiency of the SD effect. We expect our results to pave the way for switching between superconducting, metallic, and insulating behavior via an electric current.